OpenStudy (anonymous):

5x - 6 = y 2x - 3y = 4 Solve the system of equations by substitution. A. (-10/13, -128/13) B. (14/13, -8/13) C. no solution D. coincident

OpenStudy (goformit100):

You can use matrices to solve this type of questions fastest.

OpenStudy (anonymous):

we know that 5x - 6 = y so we sub 5x - 6 in for y in the other equation.. 2x - 3y = 4 2x - 3(5x - 6) = 4 2x - 15x + 18 = 4 -13x = 4 - 18 -13x = -14 x = 14/13 now sub 14/13 in for x 5x - 6 = y 5(14/13) - 6 = y 70/13 - 6 = y 70/13 - 78/13 = y - 8/13 = y answer is (14/13, -8/13)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

No i get it thank you i was just substituting the wrong thing

OpenStudy (anonymous):

glad I could help :)